# ExoMiner++ 2.0: Vetting TESS Full-Frame Image Transit Signals

### v1

@@ Abstract
The Transiting Exoplanet Survey Satellite (TESS) Full-Frame Images (FFIs) provide photometric time series for millions of stars, enabling transit searches beyond the limited set of pre-selected 2-minute targets. However, FFIs present additional challenges for transit identification and vetting. In this work, we apply ExoMiner++ 2.0, an adaptation of the ExoMiner++ framework originally developed for TESS 2-minute data, to FFI light curves. The model is used to perform large-scale planet versus non-planet classification of Threshold Crossing Events across the sectors analyzed in this study. We construct a uniform vetting catalog of all evaluated signals and assess model performance under different observing conditions. We find that ExoMiner++ 2.0 generalizes effectively to the FFI domain, providing robust discrimination between planetary signals, astrophysical false positives, and instrumental artifacts despite the limitations inherent to longer cadence data. This work extends the applicability of ExoMiner++ to the full TESS dataset and supports future population studies and follow-up prioritization.
@@ I Introduction
The Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) mission has conducted a high-precision and wide photometric survey of the sky over the past seven years and unearthed more than 7,500 planet candidates while enabling the confirmation of more than six hundred new planets. Now into its third Extended Mission (EM3), TESS continues to expand its sky survey through a series of 27-day observation cycles, each covering a $24^{\circ}\times 96^{\circ}$ field of view. The observatory collects data using mainly two different modes: 2-minute sampling of up to 20,000 targets per sector and full-frame images (FFI). [^1] Target pixel files and light curves are generated for up to 3000 20-second targets per sector since sector 27, but these are not searched for transiting planet signatures. All 20-second targets are also observed at the 2-minute cadence. For the Prime Mission, the FFI cadence was set to 30 minutes, decreasing to 10 minutes for the First Extended Mission, and to 200 seconds for the following EM2 and EM3. By reducing the cadence for FFI data, the extended missions provide increased sensitivity to short duration transits for more targets and enable the study of transient stellar events. The TESS Science Processing Operations Center (SPOC; Jenkins et al., 2016) pipeline has been used since the start of the TESS Mission to process 2-minute data to generate data products that include target pixel and light curve files for the different sectors, and the results of transiting planet searches that yield catalogs of Threshold Crossing Events (TCEs), periodic dimming events that might be consistent with planetary transits. The SPOC also produces reports and time series for TCEs that can be used towards vetting and validation efforts. For FFI data, the SPOC pipeline has been used to calibrate FFIs and assign world-coordinate system information to the FFI data delivered to the Mikulski Archive for Space Telescopes [^2] https://archive.stsci.edu/ (MAST). Starting with Year 2, the SPOC began processing 30-min targets selected from the FFIs to create calibrated target pixel files and light curves for up to 160,000 targets per sector in the TESS Northern hemisphere, and has continued to do so, with results released up to Year 6 (as of 11 December 2025, through Sector 78). Furthermore, since Sector 36 (Year 3), the SPOC has been running transiting planet searches on the hundreds of thousands of targets selected, yielding Data Validation (DV; Twicken et al., 2018; Li et al., 2019) products similar to those created for the 2-minute cadence data (for those targets with at least one detected TCE). All of these data products have been delivered to MAST as High-Level Science Products [^3] https://archive.stsci.edu/hlsp/tess-spoc/ (HLSP) for FFI targets from Sectors 36 onwards (Caldwell et al., 2020).
@@ II Related Work
Regarding the application of transit signal classification frameworks to TESS data, there are a few works that include: (1) ExoMiner++ (Valizadegan et al., 2025), a multi-branch convolutional neural network (CNN) designed to mimic subject matter experts who examine various false positive tests in the reports produced by the SPOC pipeline DV module. ExoMiner++ processes several flux time series, periodogram, difference image data, and statistical diagnostics to vet 2-minute cadence TCEs from the SPOC pipeline up to the end of Year 5; (2) AstroNet-Vetting developed for TESS FFI data (Yu et al., 2019), which is a multi-branch CNN built on the architecture of AstroNet (Shallue & Vanderburg, 2018) and adds as input a few more diagnostic tests generated by the Quick Look Pipeline (Huang et al., 2020a, b) from calibrated FFI data for early Sectors 1–6; and (3) the work in Tey et al. (2023), another modified version of AstroNet, conduct a triaging of QLP FFI TCEs using data from Years 1 through 3, with the main focus of separating periodic eclipsing signals (which include both planets and non-contact eclipsing binaries) from other astrophysical phenomena and instrumental noise.
@@ III.2 Sector Runs
The dataset of TCEs used in this work can be split into two main sets: the SPOC TCEs generated by runs using 2-minute data (from now on referred to as 2-min TCEs), and those generated by the SPOC runs on the FFI data (from now on referred to as FFI TCEs). As with any SPOC DV product, those TCE datasets are available at MAST. The following sector runs were considered for each one of these two populations:
  • 2-min TCEs: single-sector runs from Sector 1 through Sector 88, and seventeen multi-sector runs that include Sectors 1–2, 1–6, 1–9, 1–13, 1–36, 1–39, 1–46, 1–65, 1–69, 2–72, 14–19, 14–23, 14–26, 14–41, 14–50, 14–55, 14–60, 14–78, 42–43, and 42–46.
  • FFI TCEs: single-sector runs from Sector 36 through Sector 72, and one multi-sector run for Sectors 56–69.
@@ III.3 Label Sources and Assignment
We follow the procedure described in (Valizadegan et al., 2025) and use the Exoplanet Follow-up Observing Program (ExoFOP) TOI catalog [^4] https://exofop.ipac.caltech.edu/tess/ downloaded September 22, 2025, Prsa’s TESS eclipsing binary (EB) catalog (Prša et al., 2022), and TESS-ExoClass (TEC; Burke et al., 2020) tables for SPOC 2-minute and FFI TCEs (for 2-minute TCEs, we use TEC results for all single-sector runs from Sector 1 through Sector 41; for FFI TCEs, we use results from single-sector runs in range 40–73). From the ExoFOP TOI catalog, we use the confirmed planet (CP), known planet (KP), and false positive (FP) dispositioned by the TFOP Working Group (TFOPWG). We consider the KP and CP ExoFOP TOIs that are known brown dwarfs (BDs) based on TFOP Subgroup 1 (SG1) TOI catalog photometric dispositions as false positives and label them as such. TEC flux triage results are used to label TCEs as non-transiting phenomenon (NTP) and Prsa’s TESS EB catalog was used to obtain EB TCEs. At the end of this labeling process, the TCEs that lack a label are labeled as unknown (UNK). To obtain a binary label, TCEs dispositioned as CP and KP are labeled ‘planets’ and those dispositioned as BD, EB (i.e., Prsa’s EBs), FP, and NTP are labeled ‘non-planets’. Refer to Valizadegan et al. (2025) for the details of this labeling procedure.
@@ III.4 Data Preprocessing
Figure 1: From left to right, the panels display the difference, out-of-transit, and SNR fluxes for TESS SPOC 2-min TCE TIC 83053699-1-S57 (TOI 4002.01). The neighboring stars and target’s TIC-8 coordinates mapped to the CCD frame are identified as stars and as a cross, respectively. Their color encodes the stars’ $T_{mag}$.
Figure 2: Preprocessed and unnormalized neighboring stars image for TESS SPOC 2-min TCE TIC 83053699-1-S57 (TOI 4002.01). The target’s TIC-8 position is shown as a red cross.
@@ III.4.1 Encoding Neighboring Stars’ Location and Magnitude
To encode the location and magnitude of neighboring stars, we queried the TIC-8 catalog to search for stars in a search radius of $168\,\text{arcsec}$ ($\approx 8\,\text{px}$) around the target’s TIC coordinates. For each target and sector run of interest, we used the World Coordinate System (WCS) stored in the corresponding light curve file to map the celestial coordinates of the neighboring stars into pixel coordinates. Figure 1 shows an example of such search for TIC 83053699 in Sector 57. The two images – difference and out-of-transit fluxes for TCE 1 (TOI 4002.01, a false positive attributed to a nearby eclipsing binary) – show the location of the neighboring stars that were found in a search radius of $168\,\text{arcsec}$, colored as a function of their magnitude value obtained from TIC-8.
@@
  • Neighbors whose brightness indicates that even a full eclipse could not account for the observed transit depth are excluded. Let $T_{mag}^{n}$ and $T_{mag}^{t}$ be the TESS magnitude of a neighboring star and the TESS magnitude of the target, respectively, and let $\delta$ be the estimated fractional transit depth of the TCE detected for the target. A neighboring star is excluded if magnitude difference satisfies: $T_{mag}^{n}>T_{mag}^{t}-2.5*log_{10}(\delta)$. For neighbors dimmer than this threshold, a complete eclipse would be insufficient to reproduce the measured transit depth in the target light curve.
  • All neighboring sources within a pixel are excluded except for the brightest one, when multiple neighbors are present, as the brightest source is the most likely contributor to any off-target transit signal.
@@
Table 1: TESS 2-min SPOC TCE and Event Counts.
@@
After excluding those neighboring stars, we set the value of a pixel to the ratio of the target’s magnitude to the neighbors’ magnitude if there is a neighboring star; otherwise, it is set to zero. With this encoding, pixels with values greater than one represent neighbors that are brighter than the target star, while those with value less than one are dimmer. Figure 2 shows the neighbors’ image for the same TCE in Figure 1 after preprocessing the neighbors data by following these steps. Contrary to the normalization performed for the difference, out-of-transit, and SNR flux images, where we standardize the images using statistics computed from the training set, for the neighbors images we conduct fixed min-max normalization. This choice was motivated by the fact that many pixels have no neighboring stars, resulting in ratio values of zero and consequently an extremely small standard deviation of pixel values, which is required for standardization. Such small standard deviation can cause the non-zero values to become disproportionately large, potentially disrupting model training, hindering convergence, and impairing overall learning. The target-to-neighbor magnitude ratio of all targets in our dataset is not larger than 5, meaning that any neighbor’s magnitude is not more than five magnitudes larger than the target star. Because of that, we use a factor of 5 to normalize these ratios, ensuring that they are in $[0,1]$.
@@
Figure 3: Target population for the 2-min and FFI TCE datasets as a function of $T_{mag}$.
@@ III.5 Dataset
Tables 1 and 2 show the distribution of TCEs and events in the 2-min and FFI labeled datasets as a function of their binary label (i.e., planet or non-planet class) and disposition (i.e., subclass). Given that more than one TCE can be associated with the same event (e.g., TOI) due to the observation behavior of TESS, these tables also break down the dataset in terms of these unique transiting objects. NTPs are not federated to any transit event given that they are not associated with real astrophysical transiting phenomena. For that reason, we have excluded these TCEs from the event count.
As indicated by the numbers reported in these two tables, the 2-minute dataset contains a larger number of TCEs confirmed as ExoFOP TOIs than the FFI dataset. This difference is primarily due to the more extensive sector coverage, particularly the multi-sector observations, available for the 2-minute data, which increases the likelihood of detecting longer-period TOIs and planets with smaller radii, and to the smaller number of sectors with HLSP SPOC FFI data products available at this moment.
Figure 3 shows the distribution of targets in the 2-min and FFI TCE datasets (including unlabeled TCEs) as a function of $T_{mag}$. The two distributions have a significant overlap, although the FFI target population exhibits a higher concentration of targets around $T_{mag}=[9,13]$, and the 2-minute target population distribution shows a slightly heavier tail for both brighter and dimmer targets. These distinct distributions reflect the differing selection criteria for 2-min and SPOC FFI targets, highlighting the potential for discovering new planetary candidates within the remaining unvetted TCEs.
@@ IV Model
  • Replacing max pooling with strided convolution: We replaced the max-pooling layer at the end of each convolutional block with a strided convolutional layer applied at the block input (i.e., the first convolutional layer uses a stride greater than one and no padding, resulting in a reduction of the feature-map resolution). A strided convolution performs downsampling by sliding the filter over the input feature maps with a step size larger than one, thereby reducing the spatial dimensions of the output. This modification is inspired by state-of-the-art architectures, such as more recent variants of ResNet (He et al., 2016) and YOLOv5 (Redmon et al., 2016), which replace pooling operations with strided convolutions. By adopting strided convolutions, the model is able to learn task-specific downsampling operations, rather than relying on a fixed transformation such as max pooling.
@@
Table 3: Binary classification performance of ExoMiner models for the test sets of the 5-fold cross-validation experiment using multi-source data from both 2-min and FFI sources. Precision, recall, and accuracy are computed at a classification threshold of 0.5. Mean and standard deviation estimates are computed across the CV iterations. The best performer is highlighted in bold. Differences are not claimed to be statistically significant unless stated otherwise.
@@
Table 4: Subclasses recall at classification threshold of 0.5 for the test sets of the 5-fold cross-validation experiment using multi-source data from both 2-min and FFI sources. Mean and standard deviation estimates are computed across the CV iterations. The best performer is highlighted in bold. Differences are not claimed to be statistically significant unless stated otherwise.

### v2

@@ Abstract
[…]
@@ I Introduction
The Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) mission has conducted a high-precision and wide photometric survey of the sky over the past seven years and unearthed more than 7,500 planet candidates while enabling the confirmation of more than six hundred new planets. Now into its third Extended Mission (EM3), TESS continues to expand its sky survey through a series of 27-day observation cycles, each covering a $24^{\circ}\times 96^{\circ}$ field of view. The observatory collects data using mainly two different modes: 2-minute sampling of up to 20,000 targets per sector and full-frame images (FFI) [^1] Target pixel files and light curves are generated for up to 3000 20-second targets per sector since sector 27, but these are not searched for transiting planet signatures. All 20-second targets are also observed at the 2-minute cadence.. For the Prime Mission, the FFI cadence was set to 30 minutes, decreasing to 10 minutes for the First Extended Mission, and to 200 seconds for the following EM2 and EM3. By reducing the cadence for FFI data, the extended missions provide increased sensitivity to short duration transits for more targets and enabled the study of transient stellar events. The TESS Science Processing Operations Center (SPOC; Jenkins et al., 2016) pipeline has been used since the start of the TESS Mission to process 2-minute data to generate data products that include target pixel and light curve files for the different sectors, and the results of transiting planet searches that yield catalogs of Threshold Crossing Events (TCEs), periodic dimming events that might be consistent with planetary transits. The SPOC also produces reports and time series for TCEs that can be used towards vetting and validation efforts. For FFI data, the SPOC pipeline has been used to calibrate FFIs and assign world-coordinate system information to the FFI data delivered to the Mikulski Archive for Space Telescopes [^2] https://archive.stsci.edu/ (MAST). Starting with Year 2, the SPOC began processing 30-minute targets selected from the FFIs to create calibrated target pixel files and light curves for up to 160,000 targets per sector in the TESS Northern hemisphere, and has continued to do so, with results released up to Year 6 (as of 11 December 2025, through Sector 78). Furthermore, since Sector 36 (Year 3), the SPOC has been running transiting planet searches on the hundreds of thousands of targets selected, yielding Data Validation (DV; Twicken et al., 2018; Li et al., 2019) products similar to those created for the 2-minute cadence data (for those targets with at least one detected TCE). All of these data products have been delivered to MAST as High-Level Science Products [^3] https://archive.stsci.edu/hlsp/tess-spoc/ (HLSP) for FFI targets from Sectors 36 onwards (Caldwell et al., 2020).
@@ II Related Work
Regarding the application of transit signal classification frameworks to TESS data, there are a few works that include: (1) ExoMiner++ (Valizadegan et al., 2025), a multi-branch convolutional neural network (CNN) designed to mimic subject matter experts who examine various false positive tests in the reports produced by the SPOC pipeline DV module. ExoMiner++ processes several flux time series, periodogram, difference image data, and statistical diagnostics to vet 2-minute cadence TCEs from the SPOC pipeline up to the end of Year 5; (2) AstroNet-Vetting developed for TESS FFI data (Yu et al., 2019), which is a multi-branch CNN built on the architecture of AstroNet (Shallue & Vanderburg, 2018) and adds as input a few more diagnostic tests generated by the Quick Look Pipeline (Huang et al., 2020a, b) from calibrated FFI data for early Sectors 1–6; and (3) the work in Tey et al. (2023), another modified version of AstroNet, conduct a triaging of QLP FFI TCEs using data from Years 1 through 3, with the main focus of separating periodic eclipsing signals (which include both planets and non-contact eclipsing binaries) from other astrophysical phenomena and instrumental noise. To the best of our knowledge, this work is the first to incorporate the relative magnitude and location of known neighboring TIC stars, encoded explicitly as an image representation, into a machine learning vetting framework for TESS transit signals.
@@ III.2 Sector Runs
The dataset of TCEs used in this work can be split into two main sets: the SPOC TCEs generated by runs using 2-minute data (from now on referred to as 2-minute TCEs), and those generated by the SPOC runs on the FFI data (from now on referred to as FFI TCEs). As with any SPOC DV product, those TCE datasets are available at MAST. The following sector runs were considered for each one of these two populations:
  • 2-minute TCEs: single-sector runs from Sector 1 through Sector 88, and twenty multisector runs that include Sectors 1–2, 1–6, 1–9, 1–13, 1–36, 1–39, 1–46, 1–65, 1–69, 2–72, 14–19, 14–23, 14–26, 14–41, 14–50, 14–55, 14–60, 14–78, 42–43, and 42–46.
  • FFI TCEs: single-sector runs from Sector 36 through Sector 72, and one multisector run for Sectors 56–69.
@@ III.3 Label Sources and Assignment
We follow the procedure described in (Valizadegan et al., 2025) and use the Exoplanet Follow-up Observing Program (ExoFOP) TOI catalog [^4] https://exofop.ipac.caltech.edu/tess/ downloaded September 22, 2025, Prsa’s TESS eclipsing binary (EB) catalog (Prša et al., 2022), and TESS-ExoClass (TEC; Burke et al., 2020) tables for SPOC 2-minute and FFI TCEs (for 2-minute TCEs, we use TEC results for all single-sector runs from Sector 1 through Sector 41; for FFI TCEs, we use results from single-sector runs in range 40–72). From the ExoFOP TOI catalog, we use the confirmed planet (CP), known planet (KP), and false positive (FP) dispositioned by the TFOP Working Group (TFOPWG). We consider the KP and CP ExoFOP TOIs that are known brown dwarfs (BDs) based on TFOP Subgroup 1 (SG1) TOI catalog photometric dispositions as false positives and label them as such. TEC flux triage results are used to label TCEs as non-transiting phenomenon (NTP) and Prsa’s TESS EB catalog was used to obtain EB TCEs. At the end of this labeling process, the TCEs that lack a label are labeled as unknown (UNK). To obtain a binary label, TCEs dispositioned as CP and KP are labeled ‘planets’ and those dispositioned as BD, EB (i.e., Prsa’s EBs), FP, and NTP are labeled ‘non-planets’. Refer to Valizadegan et al. (2025) for the details of this labeling procedure.
@@ III.4 Data Preprocessing
Figure 1: From left to right, the panels display the difference, out-of-transit, and SNR fluxes for TESS SPOC 2-minute TCE TIC 83053699-1-S57 (TOI 4002.01). The neighboring stars and target’s TIC-8 coordinates mapped to the CCD frame are identified as stars and as a cross, respectively. Their color encodes the stars’ $T_{mag}$.
Figure 2: Preprocessed and unnormalized neighboring stars image for TESS SPOC 2-minute TCE TIC 83053699-1-S57 (TOI 4002.01). The target’s TIC-8 position is shown as a red cross.
@@ III.4.1 Encoding Neighboring Stars’ Location and Magnitude
To encode the location and magnitude of neighboring stars, we queried the TIC-8 catalog to search for stars in a search radius of $168\,\text{arcsec}$ ($\approx 8\,\text{px}$) around the target’s TIC coordinates. For each target and sector run of interest, we used the World Coordinate System (WCS) stored in the corresponding light curve file to map the celestial coordinates of the neighboring stars into pixel coordinates. Figure 1 shows an example of such search for TIC 83053699 in Sector 57. The two images, difference and out-of-transit fluxes for TCE 1 (TOI 4002.01, a false positive attributed to a nearby eclipsing binary), show the location of the neighboring stars that were found in a search radius of $168\,\text{arcsec}$, colored as a function of their magnitude value obtained from TIC-8.
@@
  • Neighbors whose brightness indicates that even a full eclipse could not account for the observed transit depth are excluded. Let $T_{mag}^{n}$ and $T_{mag}^{t}$ be the TESS magnitude of a neighboring star and the TESS magnitude of the target, respectively, and let $\delta$ be the estimated fractional transit depth of the TCE detected for the target. Under the simplifying assumption that the target star dominates the aperture flux ($F_{t}\gg F_{n}$), the maximum depth a contaminating neighbor can produce is approximated as $\delta\approx F_{n}/F_{t}$. A neighboring star is excluded if the magnitude difference satisfies: $T_{mag}^{n}-T_{mag}^{t}>-2.5*log_{10}(\delta)$. This threshold is conservative in the sense that it avoids prematurely excluding potential contaminants. It removes only those that are unequivocally too faint to produce the observed depth, while retaining any neighbor that could potentially be responsible.
  • All neighboring sources falling within the same pixel (after increasing image resolution by a factor of 5) are excluded except for the brightest one, when multiple neighbors are present, as the brightest source is the most likely contributor to any off-target transit signal.
@@
Table 1: TESS 2-minute SPOC TCE and Event Counts.
@@
After excluding those neighboring stars, we set the value of a pixel to the ratio of the target’s magnitude to the neighbors’ magnitude if there is a neighboring star; otherwise, it is set to zero. With this encoding, pixels with values greater than one represent neighbors that are brighter than the target star, while those with value less than one are dimmer. Figure 2 shows the neighbors’ image for the same TCE in Figure 1 after preprocessing the neighbors data by following these steps. Contrary to the normalization performed for the difference, out-of-transit, and SNR flux images, where we standardize the images using statistics computed from the training set, for the neighbors images we conduct fixed min-max normalization. This choice was motivated by the fact that many pixels have no neighboring stars, resulting in ratio values of zero and consequently an extremely small standard deviation of pixel values, which is required for standardization. Such small standard deviation can cause the non-zero values to become disproportionately large, potentially disrupting model training, hindering convergence, and impairing overall learning. The target-to-neighbor magnitude ratio of all targets in our dataset is not larger than 5, and so we use a factor of 5 to normalize these ratios, ensuring that they are in $[0,1]$.
@@
Figure 3: Target population for the 2-minute and FFI TCE datasets as a function of $T_{mag}$.
@@ III.5 Dataset
Tables 1 and 2 show the distribution of TCEs and events in the 2-minute and FFI labeled datasets as a function of their binary label (i.e., planet or non-planet class) and disposition (i.e., subclass). Given that more than one TCE can be associated with the same event (e.g., TOI) due to the observation behavior of TESS, these tables also break down the dataset in terms of these unique transiting objects. NTPs are not federated to any transit event given that they are not associated with real astrophysical transiting phenomena. For that reason, we have excluded these TCEs from the event count.
As indicated by the numbers reported in these two tables, the 2-minute dataset contains a larger number of TCEs confirmed as ExoFOP TOIs than the FFI dataset. This difference is primarily due to the more extensive sector coverage, particularly the multisector observations, available for the 2-minute data, which increases the likelihood of detecting longer-period TOIs and planets with smaller radii, and to the smaller number of sectors with HLSP SPOC FFI data products available at this moment.
Figure 3 shows the distribution of targets in the 2-minute and FFI TCE datasets (including unlabeled TCEs) as a function of $T_{mag}$. The two distributions have a significant overlap, although the FFI target population exhibits a higher concentration of targets around $T_{mag}=[9,13]$, and the 2-minute target population distribution shows a slightly heavier tail for both brighter and dimmer targets. These distinct distributions reflect the differing selection criteria for 2-minute and SPOC FFI targets, highlighting the potential for discovering new planetary candidates within the remaining unvetted TCEs.
@@ IV Model
  • Replacing max pooling with learned downsampling: Instead of using a maxpooling layer at the end of each convolutional block, the first convolutional layer in each block applies valid padding with a unit stride, which reduces the spatial resolution similarly to a pooling operation but allows the model to learn the downsampling transformation. This modification is inspired by state-of-the-art architectures, such as more recent variants of ResNet (He et al., 2016) and YOLOv5 (jocher2020ultralytics), which replace pooling operations with learned downsampling so the model is able to learn task-specific downsampling operations, rather than relying on a fixed transformation such as maxpooling.
@@
Table 3: Binary classification performance of ExoMiner models for the test sets of the 5-fold cross validation experiment using multi-source data from both 2-minute and FFI sources. Precision, recall, and accuracy are computed at a classification threshold of 0.5. Mean and standard deviation estimates are computed across the CV iterations. The best performer is highlighted in bold. Differences are not claimed to be statistically significant unless stated otherwise.
@@
Table 4: Subclasses recall at classification threshold of 0.5 for the test sets of the 5-fold cross validation experiment using multi-source data from both 2-minute and FFI sources. Mean and standard deviation estimates are computed across the CV iterations. The best performer is highlighted in bold. Differences are not claimed to be statistically significant unless stated otherwise.